# Bike-Bench: A Bicycle Design Benchmark for Generative Models with Objectives and Constraints

### v1

@@ 8 Conclusion and Future Directions
We proposed Bike-Bench, a constrained engineering design benchmark for generative models, comprised of 10 multiphysics and human-centered objectives and 15 geometric and safety constraints. We benchmarked a variety of design generation procedures including an LLM, tabular generative models, optimization algorithms, and optimization-augmented generative models. Our benchmarking results suggest that tabular generative models supersede other baselines in similarity performance, optimization outperforms in validity and optimality, optimization-augmented generative models reach a strong middle ground, and LLMs generally underperform.
We encourage further benchmarks of generative models, optimization algorithms, and design generation procedures that transcend boundaries. LLMs with larger context windows or tabular data adapters, Vision Language Models (VLMs), foundation models for tabular generation, and optimization-augmented mixed-variable models may be of particular interest. Lastly we advocate for the development of more engineering design benchmarks which, alongside Bike-Bench, may hopefully expand the frontier of generative AI to successful engineering design automation and beyond.
Appendix A Appendix A: Extended Details on Background, Datasets, Evaluation Criteria, Metrics, and Models
A.1 Background: Data-Driven Bicycle Design
@@
Data and computation has played an increasingly important role in bicycle design science [69], predominantly focusing on computational simulation and optimization [70, 71, 72]. However, little published research had explored bicycle design-space exploration or big-data applications before the BIKED [39] dataset was released in 2021. BIKED demonstrated some of the first applications of design-space modeling, deep learning, and generative AI in bicycle design.
@@
A.2 Datasets and evaluation criteria
A.2.1 Details on Data Collection from Human Subjects
All participants in our human subject data collection were adults (aged 18 or order), paid an hourly rate of 9 GBP. All participants were informed that their data may be used for the training and evaluation of ML/AI models and consented to this. The study passed relevant institutional review procedures without issue.
A.2.2 Building a Classification Dataset for Usability
Bike-Bench uses the binary (yes/no) ratings collected from human raters to predict the proportion of raters that would consider a design ‘easy to use.’ Rather than averaging the scores for each bike, yielding the distribution of scores shown in 3, ratings can alternatively be aggregated to form a simple classification dataset. To assess the overall consensus, factoring in a margin of error resulting from potential sampling inaccuracies, we can use a two-sided binomial test. Taking the population size of the group with the smallest number of valid users, 46, we evaluate the $99\%$ confidence interval to be approximately 20%. Thus, for any design with at least 70% consensus among respondents, we can be 99% confident in the consensus value. In other words, an average score of at least 0.7 confidently indicates that more than half of raters would consider the bike usable, whereas an average of at most 0.3 indicates that less than half of raters would consider the bike usable. Following these thresholds, 49 bikes are identified as usable and 51 as unusable, resulting in a total of 100 confident classification labels.
Figure 3: Distribution of average usability scores in Bike-Bench’s human-sourced dataset
A.2.3 Limitations and Assumptions of Datasets and Evaluators
Bike-Bench makes numerous key assumptions that may cause significant inaccuracies in design evaluation. A few key assumptions and limitations are listed below, but this is not a comprehensive list. We encourage practitioners using models trained on Bike-Bench’s data or using Bike-Bench’s design evaluation tools for real-world bicycle design to undertake additional validation and verification steps.
  • Design Representation: Bike-Bench’s design representation is expressive enough to cover a wide variety of designs adhering to a ‘conventional’ diamond-frame bike. It does not span entirely different topological layouts of bicycles, such as recumbent bikes, nor does it model complex suspension systems, such as mountain bike rear suspensions.
  • Dataset bias: The original dataet of 4500 human-design bikes features certain biases, particularly in less visually-prominent features like tube thickess values. We refer readers to BIKED [39] and FRAMED [40] for more in-depth discussion.
@@ A.2.3 Limitations and Assumptions of Datasets and Evaluators
  • Structural Evaluation: Bike-Bench assumes isotropic material properties, and substitutes any inherently non-isotropic materials (bamboo, carbon, other) with steel for the purposes of evaluation. Real metal tubes nonetheless display some anisotropic properties, causing incurabilities. Simulations are based on a simplified 3D frame model which approximates tube joints and does not incorporate certain frame features, such as wheel cutouts. Detailed discussion and comparisons to experimental validation are included in FRAMED [40]
  • Aerodynamics: The assessment of aerodynamic drag in a single speed and direct headwind and based on a basic cyclist model is a simplified assessment of bicycle aerodynamics. Real cycling undergoes a variety of different wind speeds at different angles, causing aerodynamic interactions in different flow regimes.
  • Aesthetics: Although the trained parametric-to-CLIP embedding model suffices to capture subjective details (e.g. futuristic cyberpunk-style) [76], it struggles to capture technical details of bicycle components (e.g. number of chainrings or disk versus rim brakes).
A.2.4 Visualization of geometrically infeasible bikes
@@
Figure 4: Bike-bench features numerous closed-form constraint checks to identify common geometric infeasibilities such as disconnected or colliding components (left 3). Less common infeasibilities in frame geometry are flagged if the design causes errors during 3D reconstruction (right 3).
A.3 Models and metrics
A.3.1 Details of Hypervolume Calculation
In multi-objective design optimization, designs are often compared in the objective space – the space comprised of the set of all objective scores. One design “dominates” the other if it is superior in every single objective. To calculate the hypervolume metric of a sample set, we measure the hypervolume of the region in the objective space where a corresponding design would be dominated by any design in the sample set. To keep hypervolume bounded, it is typically calculated with respect to some reference point, which we select to be the combination of maximum (worst) objective scores in every unique objective, calculated by evaluating every design in BIKED with a random conditional configuration.
A.3.2 Standardization of computation budget
Although exact comparison of computation budget across classes of methods was deemed out of scope, we made reasonable efforts to standardize computational budget allocations. Excepting the closed-source o4-mini model, which was queried through API calls, all methods were benchmarked on the same workstation (5900x processor and 3090Ti GPU) and tuned to yield maximum performance with a maximum of 1 hour for optimization, model training and inference.
A.3.3 Summary of Model Features
A summary of select features and attributes of the various design generation methods tested is compiled in Table 3.
Table 3: Select features and attributes of various design generation methods
Mixed Datatype Calls Evaluators Data-driven Conditional Generation
o4-mini ×
CTGAN × ×
TVAE × ×
OA-GAN ×
OA-VAE ×
OA-DDPM ×
OA-DDPM-G ×
Grad-Agg × × ×
EPO × × ×
NSGA-II × ×
A.3.4 Details on performance-augmented model training:
The DPP-based auxiliary loss proposed in PaDGAN [61] uses a single aggregate design quality metric. To condense bike-bench’s many objectives and constraints into a differentiable aggregate quality metric, we propose the following aggregation scheme:
@@ A.3.4 Details on performance-augmented model training:
All $n_{o}$ objectives, $o_{i}$, and $n_{c}$ constraints, $c_{i}$, are scaled by weighting parameters, $w_{o_{i}}$ and $w_{c_{i}}$. Each parameter is set automatically based on the mean absolute value of the scores seen through a random pairing of each of the $n_{D}$ points in the dataset, $D$, and randomly-sampled conditions, $C_{j}$:
@@
This continuous penalty function linearly increases for $x>0$, and morphs into an increasingly gradual decay for $x<0$, and also features a continuous first derivative. We use $\alpha=10$ and $\beta=10$ in our testing. Modulating these terms can encourage a model to focus more on objectives or constraints. For details on the DPP-based loss, given an aggregate quality function, we refer readers to [61].
A.3.5 LLM Prompts
A series of three prompts is given to the LLM model. Following the first prompt, the dataset descriptions (see Sec. A.5) and evaluation criterion descriptions (see Sec. A.4), along with a short introduction of each (see codebase) are given to the model. Following the second prompt, the condition, design, and scoring example data is given to the model. Following the third prompt, the test condition is given to the model. Conditional information is concatenated into a text string of the form: ‘Rider Body Dimensions: Upper leg length - [ULL], Lower leg length - [LLL], Arm length - [AL], Torso length - [TL], Neck and head length - [HNL], Torso width - [TW]. Use Case: [Road Biking/Mountain Biking/Commuting]. Marketing Description: [Text Prompt]’. The prompts are listed as follows:
  • I will ask you to create some bicycle designs. The bicycle designs are subject to a set of conditions: a text prompt, some rider dimensions, and a use case. Each design is defined by 70 variables, which I will describe. Some of these are categorical variables, and I will provide you with the possible values for these variables. Others are continuous. I will describe the design variables shortly. Designs are evaluated according to a set of 25 criteria. The first 10 are objectives, while the last 15 are constraints. Here are the design variable descriptions and the evaluation criteria:
  • I will also provide a dataset of existing bicycle designs. These are useful as a reference point, because it may be difficult to satisfy constraints and objectives if you deviate too far from the space of existing designs. I will also provide a set of objective scores for these bicycle designs. Since these criteria are dependent on the conditional information, each bike design is paired with a randomized condition set for the purpose of evaluation. Thus, I will provide: 1) a text file with 10 rows of conditions, 2) a csv file with 10 rows of bikes and 70 columns of parameters, and 3) a csv file of 10 sets (rows) of 25 scores (columns). Here are these files. Please have a look to try to gain an understanding of the design space and the evaluation criteria.
  • Having examined the existing designs and the evaluation of these designs alongside their associated conditions, I hope you have gained an understanding of the design space and design objectives. I will now ask you to create bicycle designs. Please deliberate on a strategy for creating high-performing designs that satisfy the constraints and objectives. I will provide you with 10 new conditions. Please create 10 unique bicycle design that satisfies the constraints and objectives. Important: You are not allowed to generate the same bike 10 times. Each design must be unique! Please provide the designs in a 10x70 csv file. No index or headers. Only the 10x70 values. Here are the 10 conditions. Remember! Every design must be unique.
A.4 Summary of Constraints and Objectives
@@
Table 4: Summary of Bike-Bench’s evaluation functions, classified as objectives or constraints. Inputs and evaluator type are also specified.
@@
Geometric Feasibility 0 13 Closed-Form + Predictor Bike
@@
Ergonomics 3 0 Closed-Form Bike + Rider + Use case

### v2

@@ 8 Conclusion and Future Directions
We proposed BikeBench, a constrained engineering design benchmark for generative models, comprising 10 multiphysics and human-centered objectives and 40 geometric, ergonomic, and safety constraints. We benchmarked a variety of design generation procedures, including an LLM, tabular generative models, optimization algorithms, and optimization-augmented generative models. Our benchmarking results suggest optimization-augmented models achieve well-rounded performance, optimization and tabular generative models achieve mixed results, and LLMs generally underperform.
We encourage further benchmarks of generative models, optimization algorithms, and design generation procedures that transcend boundaries. LLMs with larger context windows or tabular data adapters, foundation models for tabular generation, cross-conditional amortization of optimization, and optimization-augmented, mixed-variable models may be of particular interest. Lastly, we advocate for the development of more multimodal engineering design benchmarks of complex systems. Alongside BikeBench, such benchmarks will help expand the frontier of generative AI in engineering design and beyond.
Acknowledgments and Disclosure of Funding
We thank Amin Heyrani Nobari for his contributions to the renderer and Ankur Kapileshwar for his contributions to the geometric validation functions. We also acknowledge support from the MIT MISTI–Netherlands Program.
Appendix A Appendix A: Extended Results
A.1 Scorecards for Main Paper Results
We include scorecards for all models benchmarked in the full paper. First, we show unconditional generation in Figures 5 and 6. Next, we include conditional generation scorecards in Figures 7 and 8.
Figure 5: Scorecards for unconditional benchmarking results (part 1)
Figure 6: Scorecards for unconditional benchmarking results (part 2)
Figure 7: Scorecards for conditional benchmarking results (part 1)
Figure 8: Scorecards for conditional benchmarking results (part 2)
A.2 Generation Under Masked Constraints
We include benchmarking results for the masked constraint case under both conditional and unconditional generation in Table 2. GPT-5-high is excluded here due to time and budgetary constraints. In general, models with lower similarity scores struggle more to infer constraints from data. These models therefore tend to suffer a larger reduction in design quality and increase in constraint violation compared to the unmasked generation case. Few methods compete with the simple dataset subsampling baseline. This broadly suggests significant room for improvement of design synthesis algorithms in implicit constraint satisfaction.
Table 2: Design quality, constraint violation, similarity to dataset, and design diversity scores for unconditional and conditional benchmarking cases with masked constraints. Models are separated by class (LLM / Tabular-GM / OA-GM / Optimizer). The best non-baseline scores in each metric are bolded. Models are benchmarked with an evaluation budget of 1M.
Masked Generation Conditional + Masked Generation
Qual. ($\uparrow$) Viol. ($\downarrow$) Sim. ($\downarrow$) Div. ($\downarrow$) Qual. ($\uparrow$) Viol. ($\downarrow$) Sim. ($\downarrow$) Div. ($\downarrow$)
Dataset 0.0035 2.699 0.005 12.72 0.0035 2.699 0.005 12.72
Data subset 0.0151 0.123 0.082 12.41 0.0151 0.216 0.076 12.51
CTGAN 0.0079 1.046 0.123 9.94 0.0109 1.301 0.098 11.65
TVAE 0.0150 0.616 0.164 13.55 0.0151 0.428 0.154 13.54
OA-GAN 0.0128 0.505 0.196 14.02 0.0111 2.390 0.105 14.12
OA-VAE 0.0142 0.971 0.341 14.99 0.0157 1.062 0.284 14.92
OA-DDPM 0.0087 1.541 0.101 11.06 0.0007 2.495 0.201 12.93
Grad. Descent 0.0161 0.683 0.279 12.09 - - - -
NSGA-II 0.0011 3.566 0.606 11.59 - - - -
Appendix B Details on Background, Datasets, Evaluation Criteria, Metrics, and Models
B.1 Background: Data-Driven Bicycle Design
@@
Data and computation have played an increasingly important role in bicycle design science [71], predominantly focusing on computational simulation and optimization [72, 73, 74]. However, little published research had explored bicycle design-space exploration or big-data applications before the BIKED [42] dataset was released in 2021. BIKED demonstrated some of the first applications of design-space modeling, deep learning, and generative AI in bicycle design.
@@
B.2 Datasets and Evaluation Criteria
B.2.1 Details on Data Collection from Human Subjects
All participants in our human subject data collection were adults (aged 18 or older), paid an hourly rate of 9 GBP. All participants were informed that their data may be used for the training and evaluation of ML/AI models and consented to this. The study passed relevant institutional review procedures without issue.
B.2.2 Building a Classification Dataset for Usability
BikeBench uses the binary (yes/no) ratings collected from human raters to predict the proportion of raters that would consider a design ‘easy to use.’ Rather than averaging the scores for each bike, yielding the distribution of scores shown in Figure 9, ratings can alternatively be aggregated to form a simple classification dataset. To assess the overall consensus, factoring in a margin of error resulting from potential sampling inaccuracies, we can use a two-sided binomial test. Taking the population size of the group with the smallest number of valid users (46), we evaluate the $99\%$ confidence interval to be approximately 20%. Thus, for any design with at least 70% consensus among respondents, we can be 99% confident in the consensus value. In other words, an average score of at least 0.7 confidently indicates that more than half of raters would consider the bike usable, whereas an average of at most 0.3 indicates that less than half of raters would consider the bike usable. Following these thresholds, 49 bikes are identified as usable and 51 as unusable, resulting in a total of 100 confident classification labels.
Figure 9: Distribution of average usability scores in BikeBench’s human-sourced dataset.
B.2.3 Limitations and Assumptions of Datasets and Evaluators
BikeBench makes numerous key assumptions that may cause significant inaccuracies in design evaluation. A few key assumptions and limitations are listed below, but this is not a comprehensive list. We encourage practitioners using models trained on BikeBench’s data or using BikeBench’s design evaluation tools for real-world bicycle design to undertake additional validation and verification steps.
  • Design representation: BikeBench’s design representation is expressive enough to cover a wide variety of designs adhering to a ‘conventional’ diamond-frame bike. It does not span entirely different topological layouts of bicycles, such as recumbent bikes, nor does it model complex suspension systems, such as mountain bike rear suspensions.
  • Dataset bias: The original dataset of 4500 human-design bikes features certain biases, particularly in less visually-prominent features like tube thickness values. We refer readers to BIKED [42] and FRAMED [43] for more in-depth discussion.
  • Surrogate model robustness: To keep BikeBench fast and portable, we resorted to surrogate models for many of the evaluations. These may lack robustness outside of the support of their data.
@@ B.2.3 Limitations and Assumptions of Datasets and Evaluators
  • Structural evaluation: BIKED includes non-isotropic materials (bamboo, carbon, other) which are difficult to simulate without extensive assumptions. We substitute these materials with aluminum (or steel for carbon fiber) in BikeBench. Real metal tubes nonetheless display some anisotropic properties, causing inaccuracies. Simulations are based on a simplified 3D frame model which approximates tube joints and does not incorporate certain frame features, such as wheel cutouts. Detailed discussion and comparisons to experimental validation are included in FRAMED [43].
  • Aerodynamics: The assessment of aerodynamic drag in a constant direct headwind and based on a basic cyclist model is a simplified assessment of bicycle aerodynamics. Real cycling undergoes a variety of different wind speeds at different angles, causing aerodynamic interactions in different flow regimes.
  • Aesthetics: Although the trained parametric-to-CLIP embedding model suffices to capture subjective details (e.g. ‘jet-black triathlon bike’) [78], it struggles to capture technical details of bicycle components (e.g. number of chainrings or fork style).
While these inaccuracies may limit BikeBench’s robustness and utility for users interested in physically constructing the bicycles they design, their impact on BikeBench’s primary objective—benchmarking—should be minor.
B.2.4 Visualization of geometrically infeasible bikes
@@
Figure 10: BikeBench features numerous closed-form constraint checks to identify common geometric infeasibilities such as disconnected or colliding components (left 3). Less common infeasibilities in frame geometry are flagged if the design causes errors during 3D reconstruction (right 3).
B.3 Models and metrics
B.3.1 Additional Metrics and Metric Implementation Details
Details of Hypervolume Calculation:
In multi-objective design optimization, designs are often compared in the objective space—the space comprising the set of all objective scores. One design “dominates” the other if it is superior in every single objective. To calculate the hypervolume metric of a sample set, we measure the hypervolume of the region in the objective space where a corresponding design would be dominated by any design in the sample set. To keep hypervolume bounded, it is typically calculated with respect to some reference point, which we select to be the combination of maximum (worst) objective scores in every unique objective with a small additional margin. Scores are calculated by evaluating every design in BIKED with a random conditional configuration.
Novelty:
BikeBench also evaluates mean design novelty. We estimate the novelty of any given design as the Euclidean distance to nearest test-set design after scaling designs by test-set per-parameter variance.
Binary Validity:
While examining overall constraint violation is insightful, designs are ultimately invalid if they violate even a single constraint. Thus, we also evaluate the fraction of designs that simultaneously satisfy all design constraints.
The public leaderboard will feature these extended scores, in addition to the primary four metrics discussed in the main paper.
B.3.2 Details on optimization-augmented model training
This section presents a detailed discussion of the optimization-augmented generative model training process, with a visual overview in Figure 11.
Figure 11: Overview of the optimization-augmented generative model training process.
The DPP-based auxiliary loss proposed in PaDGAN [64] uses a single aggregate design quality metric. To condense BikeBench’s many objectives and constraints into a differentiable aggregate quality metric, we propose the following aggregation scheme:
@@ B.3.2 Details on optimization-augmented model training
All $n_{o}$ objectives, $o_{i}$, and $n_{c}$ constraints, $c_{i}$, are scaled by weighting parameters, $w_{o_{i}}$ and $w_{c_{i}}$. Each parameter is set automatically based on the mean absolute value of the scores observed from random pairings of the $n_{D}$ points in the dataset, $D$, with randomly sampled conditions, $C_{j}$:
@@
This continuous penalty function linearly increases for $x>0$, and morphs into an increasingly gradual decay for $x<0$, and also features a continuous first derivative. We use $\alpha=100$ and $\beta=10$ in our benchmarking. Modulating these terms can encourage a model to focus more on objectives or constraints. For details on the DPP-based loss, given an aggregate quality function, we refer readers to [64].
In the unconditional benchmarks, a small portion of the evaluation budget is held out during training to perform a final rejection sampling pass on a few hundred designs. This (generally) ensures a perfectly valid set. Due to relatively high raw validity rates, relatively few designs have to be passed through the rejection sampling step, making its cost very small (under 10% of the total evaluation budget).
The guided DDPM is only applied to the unconditional case, as it directly calls the evaluators on the test-set conditions. The evaluation budget is not high enough to invoke evaluator-based guidance at every denoising step. Instead, the guidance is invoked in a scheduled manner at only certain time steps.
We refer users interested in detailed training parameters for the OA-GMs to the repository linked in the abstract.
B.3.3 LLM Prompts
A series of prompts is given to the LLM model. These prompts are listed below, excepting a few of the longer subcomponents which are substantiated later in the appendix.
Full-Stack LLM Prompts: (System Prompt) “You are a large language model who will be assisting with a computational bicycle design task. The user will give you some examples of bicycle designs then ask you to create new ones. This is a challenging design task because the designs are represented using csv tables. You’$l$ have to keep good track of the order of the bicycle parameters to make sure you dont forget any or mix any up. The user will also give you some performance metrics for the existing bicycle designs. You will have to use your knowledge of physics and design principles to try to ensure that the designs you generate are valid. Importantly, the user does not want any sort of feedback. They only want the raw final designs in csv format. You’$l$ task is to figure out how to design them to try to meet the design constraints and give them cleanly to the user in csv format. ” “I will ask you to create some bicycle designs. The bicycle designs are subject to a set of conditions: a text prompt, some rider dimensions, and a use case. Each design is defined by 64 variables, which I will describe. Some of these are categorical variables, and I will provide you with the possible values for these variables. Others are continuous or boolean. I will describe the design variables shortly. Designs are evaluated according to a set of 50 criteria. 10 of these are objectives and 40 are constraints. Here are the descriptions of the design variables and the evaluation criteria:” Dataset descriptions (see Sec. B.5). Evaluation criterion descriptions (see Sec. B.4). “Before asking you to design the bikes, I will provide a dataset of existing bicycle designs to reference. These are useful as a reference point, because it may be difficult to satisfy constraints and objectives if you deviate too far from the space of existing designs. I will also provide a set of objective scores for these bicycle designs. Here are the 25 conditional inputs for the bicycle generation task:” Conditional inputs (see Sec. B.6). “Here is a csv file with 25 examples of bikes that satisfy the constraints. The csv has 25 rows (bikes) plus a header row and exactly 64 columns (parameters):” CSV file of valid designs “Next, here are the corresponding scores for those 25 bikes in a csv file of 25 designs (rows) plus a header row by 50 scores (columns). You can see that all constraint scores are less or equal to zero:” CSV file of valid designs’ scores “Now that you have seen some example designs that satisfy the constraints, as well as their scores, here is a csv file with 25 examples of bikes that violate the constraints. This csv also has 25 rows (bikes) plus a header row and exactly 64 columns (parameters):” CSV file of invalid designs “Finally, here are the corresponding scores for those 25 invalid bikes in a csv file of 25 designs (rows) plus a header row by 50 scores (columns). You can see that at least one constraint score is positive for each:” CSV file of invalid designs’ scores “Having examined 25 valid and 25 invalid designs and the evaluation of these designs, I hope you have gained an understanding of the design space and design objectives. Please deliberate on a strategy for creating high-performing designs that satisfy the constraints and objectives. Your goal is to create designs that will minimize the objective scores while ensuring every constraint score is less or equal to zero. Here is the prompt for which you will be generating designs:” String with the test-set condition “You must create 100 unique bicycle design that satisfy the constraints and objectives. Important: You are not allowed to generate the same bike 100 times. Each design must be unique! You are also encouraged to improve over the valid designs given to you, which may not be very optimal. Please provide the designs in a 100 x 64 csv file with one extra column for the index and one extra row for the column hearers. Please count up from 1-100 in the indices. For the column headers, we recommend you use the parameter names described to help you keep track of what parameter is what. Remember! Every design must be unique. Your final output must be only the csv itself, nothing more. Generate the csv file now.”
B.4 Summary of Constraints and Objectives
This section summarizes BikeBench’s 50 unique constraints and objectives. An overview is presented in Table 3, which includes the types of evaluators and the inputs to the evaluation functions.
@@ B.4 Summary of Constraints and Objectives
Table 3: Summary of BikeBench’s evaluation functions, classified as objectives or constraints. Inputs and evaluator type are also specified.
@@
Geometric Feasibility 0 32 Closed-Form + Predictor Bike
@@
Ergonomics 3 6 Closed-Form Bike + Rider + Use case